a wheel makes 6 1/12 revolutions per minute. if it rotate for 80 minutes how many revolutions does it make
Answer by edjones(8007)

6 1/12 = 73/12
73/12 * 80 = 1460/3
=486 2/3 revolutions.
.
Ed